# SKADA-Bench: Benchmarking Unsupervised Domain Adaptation Methods with Realistic Validation

### v1

@@ 4.2 Study of validation scorers
We now investigate the performance of the various scorers to select hyperparameters of the DA method. First, we consider the relationship between the cross-val score and the accuracy for each inner split. In Figure 3, we plot for each scorer the cross-val score as a function of the accuracy computed on the test set and report the Pearson correlation coefficient $\rho$. As expected, the supervised scorer is highly correlated with the accuracy ($\rho=0.98$), as it has access to the target labels. We observe that SND, DEV, and PE do not provide a good proxy to select hyperparameters that give the best-performing models ($\rho\leq 0.06$). On the contrary, IW and CircV are correlated with the accuracy, $\rho=0.53$ and $\rho=0.71$ respectively. This is coherent with their selection as the best scorer in most scenarios in Table 2. Still, while those scorers are well correlated with the target accuracy, it is important to note that they have a large variance. For instance, a score close to 1 in IW or CircV corresponds to an accuracy between 0.5 and 1.0.
Furthermore, we provide in Figures D.3 and D.4, from Appendix D, several visualizations that illustrate the relationship between the accuracy achieved when using a supervised scorer and the accuracy obtained when using different unsupervised scorers. We also visualize in Figure D.2 the drop in performance when using the best-unsupervised scorer instead of the supervised scorer. Interestingly some methods such as KMM, EntOT, and ClassRegOT can lose up to 10% accuracy when using realistic scorers, which might come from their high number of parameters or their sensitivity to them.
@@ 5 Conclusion
In this work, we introduced SKADA-Bench, a comprehensive benchmark for unsupervised domain adaptation, focusing on shallow DA methods and the impact of their model selection. Our findings reveal that few DA methods consistently perform well across diverse datasets and that model selection scorers significantly influence their effectiveness. For each DA method, we provide the optimal model selection scorer for unsupervised hyperparameter tuning based on our experiments. We believe this benchmark will aid the community in understanding the critical role of model selection in DA performance and inspire the development of methods that are more robust to hyperparameter choices.
Acknowledgments
This work was supported by grant ANR-22-PESN-0012 to AC under the France 2030 program, ANR-20-CHIA-0016 and ANR-20-IADJ-0002 to AG, and ANR-23-ERCC-0006 to YL, TG, and RF, all managed by the Agence Nationale de la Recherche (ANR). Additionally, this project received funding from the European Union’s Horizon Europe research and innovation program under grant agreement 101120237 (ELIAS).
All the datasets used for this work were accessed and processed on the Inria and IP Paris (IDCS) compute infrastructures.
@@ Acknowledgments
The entire code and results of SKADA-Bench are open-sourced at https://github.com/scikit-adaptation/skada-bench. The implementation of the DA methods and scorers is provided along with access to the simulated and real-world datasets. All the performance tables and figures can be reproduced effortlessly, and guidelines with minimal working examples are given to add new DA methods and datasets.
@@
In this appendix, we provide additional information regarding the validation procedure used in the literature for each DA method implemented in SKADA-Bench in Section A. We provide a detailed description of the data and preprocessing used in SKADA-Bench in Section B. In Section C, we give minimal working Python examples to add a new DA method and dataset in SKADA-Bench. Finally, we provide the detailed benchmark results in Section D. In particular, the results per dataset can be found in Section D.1. We discuss in Section D.2 the impact of the choice of base estimator on the performance of DA methods for the simulated datasets. The results of each DA method with the supervised scorer on all the datasets are given in Table D.12 of Section D.3, which parallels Table 2. A thorough analysis of the effect of using realistic unsupervised scorers is also provided in Section D.4. Finally, the computational efficiency of each DA method is studied in Section D.5 and the hyperparameters used for grid search are given in Section D.6. We display the corresponding table of contents below.
Appendix A Model selection in Domain Adaptation
Table A.1: Validation procedure in Domain Adaptation methods. NA stands for not applicable and means that there are no hyperparameters. None means that no validation procedure has been conducted or that it is not specified in the original paper.
@@ Appendix A Model selection in Domain Adaptation
Reweighting Density Reweight [50] None Bandwidth fixed by Silverman method
Discriminative Reweight [46] NA No hyperparameters
Gaussian Reweight [46] None Not specified in [46]
KLIEP [51] Integrated CV Likelihood CV [51] on target
KMM [21] None Fixed data-dependent hyperparameters
NN Reweight [30] None Number of neighbors fixed to one
MMDTarS [60] CV Not specified if done on source or target
Mapping Coral [52] NA No hyperparameters
OT mapping [7] CV target/CircCV Unclear in the text
Lin. OT mapping [13] NA No hyperparameters
MMD-LS [60] CV Not specified if done on source or target
Subsp. SA [12] 2-fold CV on source -
TCA [39] Validation on target Target subset used to tune parameters
TSL [47] None Not specified in [47]
Other JDOT [6] Reverse CV [61] -
OT label prop [48] NA No hyperparameters
DASVM [3] Circular Validation [3] -
In Table A.1, we provide additional information on the validation procedures used in the original papers that proposed the different domain adaptation methods implemented in SKADA-Bench. The first column is the name of the method, the second column contains the procedure used to select hyperparameters and the last column provides additional details. What is striking is that many methods do not conduct or specify a validation procedure to select the hyperparameters, which limits the performance of the proposed method on a novel dataset. Several others rely on cross-validation using target data. However, since target labels are typically unavailable in practical scenarios, this validation approach is unrealistic. Overall, many methods have been evaluated with unrealistic or not reproducible validation procedures, making the performance of the proposed methods appear over-optimistic. A key contribution of our work is the extensive comparison of realistic, unsupervised scorers for selecting optimal hyperparameters and base estimators in DA methods.
Appendix B Datasets description and preprocessing
@@
First, three computer vision datasets are proposed: Office31 [26], Office Home [55], and MNIST/USPS [28]. We create embeddings for Office31 using the Decaff preprocessing method [10] and for Office Home using a pre-trained ResNet50 [19]. These embeddings, as well as vectorized MNIST/USPS, are dimensionally reduced with a Principal Component Analysis (PCA). These three datasets encompass 3, 4, and 3 domains, respectively and all pairs of adaptations are used as DA problems. MNIST/USPS contain clear and blurry images digits, Office31 differentiates between images captured by various devices, while for OfficeHome, its by image style.
@@
The second task is Natural Language Processing (NLP). Two datasets are studied: 20Newsgroup [27] and Amazon Review [33]. The 20Newsgroup dataset contains 20.000 documents categorized into 4 categories: talk, rec, comp, and sci. The learning task is to classify documents across categories. First, the documents are embedded using a Large Language Model (LLM) [43, 57], and then PCA is applied for dimensionality reduction.
@@
We propose two tabular datasets. The first one is the Mushroom dataset [8], where the task is to classify whether a mushroom is poisonous or not. The two domains are separated according to the mushroom’s stalk shape (enlarging vs. tapering). The tabular data are one-hot-encoded to transform categorical data into numerical data. The second dataset is Phishing [34]. The classification problem involves determining whether a webpage is a phishing or a legitimate one. The domains are separated according to the availability of the IP address. Since the data are already numerical, no preprocessing is done on this dataset.
@@
The last task is BCI Motor Imagery. The dataset used is BCI Competition IV [54], often used in the literature [1]. The task is to classify four kinds of motor imagery (right hand, left hand, feet, and tongue) from EEG data. In this dataset, nine subjects are available. The domains are separated based on session number. For each subject, session 1 is considered as the source domain and session 2 is considered as the target domain. The data are multivariate signals. To embed the data, we first compute the covariance and then project this covariance on the Tangent Space as proposed in [1].
Appendix C Adding new methods and datasets to SKADA-Bench
Using the benchopt framework for this benchmark allows users to easily add novel domain adaptation (DA) methods and datasets. To that end, users should adhere to the benchopt [35] conventions. We provide below the guidelines with examples in Python to add a new DA method and a new dataset to SKADA-Bench.
C.1 Adding a new DA method
@@
  • This class should implement a get_estimator() function, which returns a class inheriting from sklearn.BaseEstimator and accepts sample_weight as fit parameter. In the benchmark we used the Domain Adaptation toolbox SKADA [17] that provides many esDA estimatos with correct interface.
@@
def __init__(self, param1=10, param2='auto'):
@@
'param1': [10, 100],
'param2': ['auto', 'manual']
@@ C.2 Adding a new dataset
# sample_domain is negative for target sampels and positive for source
@@
Appendix D Benchmark detailed results
D.1 Results per datasets
@@
These detailed tables where cell in green denote a gain wrt Train Src (average outside of standard deviation of Train Src) better illustrate the challenges of domain adaptation (DA) methods. They show that not all shifts are equivalent within a given dataset. For example, Table D.5 reveals that only 4 shifts in the AmazonReview dataset present a DA problem (defined as a $>3\%$ difference in accuracy between Train Src and Train Tgt). While for the other shifts, we achieve similar performance whether we train on source or target data. Additionally, some specific shifts present a DA problem that no method can successfully address. This can be seen in the dsl → amz shift in the Office31 dataset, as shown in Table D.2. Finally, some DA methods perform consistently across all shifts within a dataset, as demonstrated by the results for the 20Newsgroup dataset in Table D.4.
Table D.1: Accuracy score for MNIST/USPS dataset for each shift compared for all the methods. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
@@
Reweighting Dens. RW 0.66 $\pm$ 0.02 0.42 $\pm$ 0.02 0.54 $\pm$ 0.02 12.75
Disc. RW 0.6 $\pm$ 0.01 0.4 $\pm$ 0.02 0.5 $\pm$ 0.02 18.00
Gauss. RW 0.11 $\pm$ 0.01 0.11 $\pm$ 0.01 0.11 $\pm$ 0.01 20.50
KLIEP 0.66 $\pm$ 0.02 0.43 $\pm$ 0.02 0.54 $\pm$ 0.02 12.50
KMM 0.64 $\pm$ 0.02 0.41 $\pm$ 0.03 0.52 $\pm$ 0.02 17.00
NN RW 0.66 $\pm$ 0.02 0.42 $\pm$ 0.02 0.54 $\pm$ 0.02 11.50
MMDTarS 0.66 $\pm$ 0.02 0.42 $\pm$ 0.02 0.54 $\pm$ 0.02 11.25
@@
MMD-LS 0.63 $\pm$ 0.05 0.41 $\pm$ 0.04 0.52 $\pm$ 0.05 15.50
Subspace JPCA 0.66 $\pm$ 0.02 0.37 $\pm$ 0.02 0.51 $\pm$ 0.02 12.50
SA 0.71 $\pm$ 0.03 0.41 $\pm$ 0.02 0.56 $\pm$ 0.02 10.50
TCA 0.1 $\pm$ 0.05 0.12 $\pm$ 0.03 0.11 $\pm$ 0.04 20.50
TSL 0.31 $\pm$ 0.05 0.19 $\pm$ 0.01 0.25 $\pm$ 0.03 19.00
Other JDOT 0.73 $\pm$ 0.02 0.53 $\pm$ 0.02 0.63 $\pm$ 0.02 3.00
@@
Table D.2: Accuracy score for Office31 dataset for each shift compared for all the methods. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
@@
Train Src 0.55 $\pm$ 0.04 0.51 $\pm$ 0.04 0.5 $\pm$ 0.02 0.9 $\pm$ 0.02 0.49 $\pm$ 0.01 0.59 $\pm$ 0.03 9.4
Train Tgt 0.93 $\pm$ 0.03 0.95 $\pm$ 0.01 0.78 $\pm$ 0.01 0.95 $\pm$ 0.02 0.77 $\pm$ 0.01 0.88 $\pm$ 0.02 1.0
Reweighting Dens. RW 0.49 $\pm$ 0.05 0.51 $\pm$ 0.04 0.5 $\pm$ 0.03 0.85 $\pm$ 0.09 0.48 $\pm$ 0.01 0.57 $\pm$ 0.04 11.6
Disc. RW 0.57 $\pm$ 0.07 0.51 $\pm$ 0.04 0.49 $\pm$ 0.03 0.87 $\pm$ 0.04 0.46 $\pm$ 0.0 0.58 $\pm$ 0.03 12.6
Gauss. RW 0.34 $\pm$ 0.04 0.1 $\pm$ 0.03 0.07 $\pm$ 0.01 0.38 $\pm$ 0.03 0.1 $\pm$ 0.01 0.2 $\pm$ 0.03 20.0
KLIEP 0.55 $\pm$ 0.04 0.51 $\pm$ 0.04 0.49 $\pm$ 0.02 0.9 $\pm$ 0.02 0.49 $\pm$ 0.01 0.59 $\pm$ 0.02 10.8
KMM 0.56 $\pm$ 0.06 0.54 $\pm$ 0.03 0.49 $\pm$ 0.03 0.86 $\pm$ 0.02 0.47 $\pm$ 0.02 0.58 $\pm$ 0.03 11.7
NN RW 0.53 $\pm$ 0.04 0.5 $\pm$ 0.03 0.49 $\pm$ 0.02 0.91 $\pm$ 0.02 0.48 $\pm$ 0.01 0.58 $\pm$ 0.02 12.7
MMDTarS 0.54 $\pm$ 0.05 0.51 $\pm$ 0.04 0.48 $\pm$ 0.02 0.76 $\pm$ 0.14 0.48 $\pm$ 0.01 0.56 $\pm$ 0.05 13.0
Mapping CORAL 0.53 $\pm$ 0.04 0.53 $\pm$ 0.03 0.5 $\pm$ 0.02 0.91 $\pm$ 0.03 0.5 $\pm$ 0.02 0.59 $\pm$ 0.03 6.6
MapOT 0.46 $\pm$ 0.04 0.52 $\pm$ 0.03 0.47 $\pm$ 0.02 0.84 $\pm$ 0.03 0.43 $\pm$ 0.02 0.55 $\pm$ 0.03 12.0
EntOT 0.52 $\pm$ 0.02 0.56 $\pm$ 0.02 0.49 $\pm$ 0.03 0.88 $\pm$ 0.03 0.47 $\pm$ 0.01 0.58 $\pm$ 0.02 12.5
ClassRegOT 0.59 $\pm$ 0.08 0.62 $\pm$ 0.03 0.49 $\pm$ 0.03 0.86 $\pm$ 0.04 0.49 $\pm$ 0.01 0.61 $\pm$ 0.04 7.0
LinOT 0.54 $\pm$ 0.05 0.55 $\pm$ 0.04 0.5 $\pm$ 0.02 0.9 $\pm$ 0.02 0.49 $\pm$ 0.01 0.59 $\pm$ 0.03 8.0
MMD-LS 0.46 $\pm$ 0.07 0.49 $\pm$ 0.06 0.49 $\pm$ 0.03 0.89 $\pm$ 0.04 0.43 $\pm$ 0.04 0.55 $\pm$ 0.05 14.7
Subspace JPCA 0.56 $\pm$ 0.03 0.51 $\pm$ 0.04 0.42 $\pm$ 0.02 0.84 $\pm$ 0.03 0.44 $\pm$ 0.02 0.55 $\pm$ 0.03 11.6
SA 0.56 $\pm$ 0.03 0.51 $\pm$ 0.03 0.49 $\pm$ 0.02 0.89 $\pm$ 0.01 0.49 $\pm$ 0.01 0.59 $\pm$ 0.02 11.3
TCA 0.04 $\pm$ 0.02 0.03 $\pm$ 0.01 0.04 $\pm$ 0.0 0.04 $\pm$ 0.0 0.04 $\pm$ 0.0 0.04 $\pm$ 0.0 21.0
TSL 0.55 $\pm$ 0.04 0.51 $\pm$ 0.03 0.5 $\pm$ 0.02 0.9 $\pm$ 0.02 0.49 $\pm$ 0.01 0.59 $\pm$ 0.03 7.4
Other JDOT 0.63 $\pm$ 0.04 0.64 $\pm$ 0.01 0.5 $\pm$ 0.02 0.77 $\pm$ 0.03 0.49 $\pm$ 0.01 0.6 $\pm$ 0.02 4.2
OTLabelProp 0.58 $\pm$ 0.03 0.63 $\pm$ 0.01 0.48 $\pm$ 0.04 0.88 $\pm$ 0.02 0.46 $\pm$ 0.01 0.61 $\pm$ 0.02 11.2
Table D.3: Accuracy score for OfficeHome dataset for each shift compared for all the methods. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.

### v2

@@ 4.2 Study of validation scorers
We now investigate the performance of the various scorers to select hyperparameters of the DA method. First, we consider the relationship between the cross-val score and the accuracy for each inner split. In Figure 3, we plot for each scorer the cross-val score as a function of the accuracy computed on the test set and report the Pearson correlation coefficient $\rho$. As expected, the supervised scorer is highly correlated with the accuracy ($\rho=0.98$), as it has access to the target labels. We observe that SND, DEV, and PE do not provide a good proxy to select hyperparameters that give the best-performing models ($\rho\leq 0.06$). On the contrary, MixVal, IW and CircV are correlated with the accuracy, $\rho=0.34$, $\rho=0.56$ and $\rho=0.71$ respectively. This is coherent with their selection as the best scorer in most scenarios in Table 2. Still, while those scorers are well correlated with the target accuracy, it is important to note that they have a large variance. For instance, a score close to 1 in IW or CircV corresponds to an accuracy between 0.5 and 1.0.
Furthermore, we provide in Figures 8 and 9, from Appendix E, several visualizations that illustrate the relationship between the accuracy achieved when using a supervised scorer and the accuracy obtained when using different unsupervised scorers. We also visualize in Figure 7 the drop in performance when using the best-unsupervised scorer instead of the supervised scorer. Interestingly some methods such as KMM, EntOT, and ClassRegOT can lose up to 10% accuracy when using realistic scorers, which might come from their higher number of parameters or their sensitivity to them.
@@
4.3 Deep DA methods
Table 3: Accuracy scores for Deep methods on selected real-life datasets using DA scorers. LinOT is reported as the overall top-performing shallow method. Green indicates that the score improved with the DA methods. The darker the color, the more significant the change.
MNIST/USPS Office31 OfficeHome BCI Selected Scorer Rank
Train Src 0.85 0.77 0.58 0.54 6.19
Train Tgt 0.98 0.96 0.83 0.56 2.07
DeepCORAL (Sun & Saenko, 2016) 0.93 0.77 0.59 0.54 MixVal 3.29
DAN (Long et al., 2015b) 0.86 0.75 0.56 0.53 IW 4.76
DANN (Ganin et al., 2016a) 0.9 0.79 0.59 0.41 MixVal 4.98
DeepJDOT (Damodaran et al., 2018b) 0.9 0.82 0.62 0.54 PE 2.92
MCC (Jin et al., 2020) 0.93 0.83 0.66 0.53 MixVal 2.38
MDD (Zhang et al., 2019) 0.87 0.78 0.56 0.4 MixVal 4.96
SPA (Xiao et al., 2023b) 0.91 0.78 0.56 0.41 DEV 5.39
LinOT (Flamary et al., 2020) 0.64 0.6 0.57 0.61 CircV
Although most of the recent work on domain adaptation focus on Deep methods for computer vision tasks, shallow methods are competitive in many applications such as tabular data (Grinsztajn et al., 2022) or datasets with a relatively small number of training examples such as BCI (Chevallier et al., 2024). Moreover, shallow methods can also benefit from recent advances in Deep learning by using Deep pre-trained feature extraction (transfer learning). However, to the best of our knowledge, the literature lacks quantitative comparison between shallow methods applied on Deep pre-trained feature extraction and Deep DA methods. To this end, we ran a benchmark using the same pipeline as in Table 2 with three Domain Invariant Deep DA methods on the CV and BCI datasets.
The results are available in Table 3 with a comparison to the best performing shallow method from Table 2. One of the most notable and expected difference is on MNIST/USPS. Shallow methods struggle to achieve good performances, even on Train Tgt, as they rely on PCA for feature extraction. Deep methods, on the other hand, use CNNs, leading to large accuracy gains on train on Src and Tgt but also on Deep DA methods. However, it is important to note that while DeepJDOT, DANN and MCC improve performance on all datasets, they remain far from the train on Tgt accuracies, partly due to the difficulty in tuning their parameters (see Appendix E.3 with the supervised scorer). The superior performance of Deep DA methods on CV datasets can be attributed to the relationship between classification in the DA subspace and the disentanglement of semantic (discriminant) content from style (domain shift) (Gonzalez-Garcia et al., 2018; Gabbay et al., 2021). Numerous studies have demonstrated that semantic embeddings can be effectively recovered, supporting the assumption that a (nonlinear) subspace shift is reasonable for CV tasks. However, for the BCI dataset, where the amount of data is limited, the performances of Deep DA methods are inferior to some other shallow methods (i.e., LinOT for example). Finally, a method like DANN, which is often considered as a baseline in the community, has been shown to be hard to validate and requires setup that can be difficult to determine across different settings. These results emphasize, that while Deep invariant DA methods can be effective, they do not consistently yield good results across modalities, whereas shallow DA methods can achieve similar or superior performances with less effort and fewer computational resources in low data regimes.
@@ 5 Conclusion
In this work, we introduced SKADA-Bench, a extensive benchmark for unsupervised domain adaptation, carefully evaluating the impact of the model selection criteria and covering diverse modalities: computer vision, natural language processing, tabular data and biosignals. While being quite comprehenvise on shallow methods, our results also provide a comparison of three common deep DA baselines on computer vision and biosignals. Importantly SKADA-Bench can be easily extended with new datasets and methods to push further the state-of-the-art. Our findings reveal that few shallow DA methods consistently perform well across diverse datasets and that model selection scorers significantly influence their effectiveness. While deep DA methods show similar trends, they often require more extensive hyperparameter tuning and architectures tailored to each modality. Notably, they tend to perform significantly better than shallow methods on some modalities, such as computer vision, while facing challenges on others such as biosignals. For each DA method, we provide the optimal model selection scorer for unsupervised hyperparameter tuning based on our experiments.
Appendix A Appendix
@@
The entire code and results of SKADA-Bench will be open-sourced and available online. The implementation of the DA methods and scorers is provided along with access to the simulated and real-world datasets. All the performance tables and figures can be reproduced effortlessly, and guidelines with minimal working examples are given to add new DA methods and datasets.
@@
In this appendix, we provide additional information regarding the validation procedure used in the literature for each DA method implemented in SKADA-Bench in Section B. We provide a detailed description of the data and preprocessing used in SKADA-Bench in Section C. In Section D, we give minimal working Python examples to add a new DA method and dataset in SKADA-Bench. Finally, we provide the detailed benchmark results in Section E. In particular, the results per dataset can be found in Section E.1. We discuss in Section E.2 the impact of the choice of base estimator on the performance of DA methods for the simulated datasets. The results of each DA method with the supervised scorer on all the datasets are given in Table 20 of Section E.3, which parallels Table 2. A thorough analysis of the effect of using realistic unsupervised scorers is also provided in Section E.6. Finally, the computational efficiency of each DA method is studied in Section E.7 and the hyperparameters used for grid search are given in Section E.8. We display the corresponding table of contents below.
Appendix B Model selection in Domain Adaptation
Table 4: Validation procedure in Domain Adaptation methods. NA stands for not applicable and means that there are no hyperparameters. None means that no validation procedure has been conducted or that it is not specified in the original paper.
@@ Appendix B Model selection in Domain Adaptation
Reweighting Density Reweight  (Sugiyama & Müller, 2005) None Bandwidth fixed by Silverman method
Discriminative Reweight  (Shimodaira, 2000) NA No hyperparameters
Gaussian Reweight  (Shimodaira, 2000) None Not specified in  (Shimodaira, 2000)
KLIEP  (Sugiyama et al., 2007b) Integrated CV Likelihood CV (Sugiyama et al., 2007b) on target
KMM  (Huang et al., 2006) None Fixed data-dependent hyperparameters
NN Reweight  (Loog, 2012) None Number of neighbors fixed to one
MMDTarS  (Zhang et al., 2013) CV Not specified if done on source or target
Mapping Coral  (Sun et al., 2017) NA No hyperparameters
OT mapping  (Courty et al., 2017b) CV target/CircCV Unclear in the text
Lin. OT mapping  (Flamary et al., 2020) NA No hyperparameters
MMD-LS  (Zhang et al., 2013) CV Not specified if done on source or target
Subsp. SA  (Fernando et al., 2013) 2-fold CV on source -
TCA  (Pan et al., 2011) Validation on target Target subset used to tune parameters
TSL  (Si et al., 2010) None Not specified in (Si et al., 2010)
Other JDOT  (Courty et al., 2017a) Reverse CV  (Zhong et al., 2010) -
OT label prop  (Solomon et al., 2014) NA No hyperparameters
DASVM  (Bruzzone & Marconcini, 2010a) Circular Validation  (Bruzzone & Marconcini, 2010a) -
In Table 4, we provide additional information on the validation procedures used in the original papers that proposed the different domain adaptation methods implemented in SKADA-Bench. The first column is the name of the method, the second column contains the procedure used to select hyperparameters and the last column provides additional details. What is striking is that many methods do not conduct or specify a validation procedure to select the hyperparameters, which limits the performance of the proposed method on a novel dataset. Several others rely on cross-validation using target data. However, since target labels are typically unavailable in practical scenarios, this validation approach is unrealistic. Overall, many methods have been evaluated with unrealistic or not reproducible validation procedures, making the performance of the proposed methods appear over-optimistic. A key contribution of our work is the extensive comparison of realistic, unsupervised scorers for selecting optimal hyperparameters and base estimators in DA methods.
Appendix C Datasets description and preprocessing
@@
First, three computer vision datasets are proposed: Office31 (Koniusz et al., 2017), Office Home (Venkateswara et al., 2017), and MNIST/USPS (Liao & Carneiro, 2015). We create embeddings for Office31 using the Decaff preprocessing method (Donahue et al., 2014) and for Office Home using a pre-trained ResNet50 (He et al., 2016). These embeddings, as well as vectorized MNIST/USPS, are dimensionally reduced with a Principal Component Analysis (PCA). These three datasets encompass 3, 4, and 3 domains, respectively and all pairs of adaptations are used as DA problems. MNIST/USPS contain clear and blurry images digits, Office31 differentiates between images captured by various devices, while for OfficeHome, its by image style.
@@
The second task is Natural Language Processing (NLP). Two datasets are studied: 20Newsgroup (Lang, 1995) and Amazon Review (McAuley et al., 2015). The 20Newsgroup dataset contains 20.000 documents categorized into 4 categories: talk, rec, comp, and sci. The learning task is to classify documents across categories. First, the documents are embedded using a Large Language Model (LLM) (Reimers & Gurevych, 2019; Xiao et al., 2023a), and then PCA is applied for dimensionality reduction.
@@
We propose two tabular datasets. The first one is the Mushroom dataset (Dai et al., 2007), where the task is to classify whether a mushroom is poisonous or not. The two domains are separated according to the mushroom’s stalk shape (enlarging vs. tapering). The tabular data are one-hot-encoded to transform categorical data into numerical data. The second dataset is Phishing (Mohammad et al., 2012). The classification problem involves determining whether a webpage is a phishing or a legitimate one. The domains are separated according to the availability of the IP address. Since the data are already numerical, no preprocessing is done on this dataset.
@@
The last task is BCI Motor Imagery. The dataset used is BCI Competition IV (Tangermann et al., 2012), often used in the literature (Barachant et al., 2012) and availaoble in MOABB (Aristimunha et al., 2023). The task is to classify four kinds of motor imagery (right hand, left hand, feet, and tongue) from EEG data. In this dataset, nine subjects are available. The domains are separated based on session number. For each subject, session 1 is considered as the source domain and session 2 is considered as the target domain. The data are multivariate signals. To embed the data, we first compute the covariance and then project this covariance on the Tangent Space as proposed in Barachant et al. (2012).
Appendix D Adding new methods and datasets to SKADA-Bench
Using the benchopt framework for this benchmark allows users to easily add novel domain adaptation (DA) methods and datasets. To that end, users should adhere to the benchopt (Moreau et al., 2022) conventions. We provide below the guidelines with examples in Python to add a new DA method and a new dataset to SKADA-Bench.
D.1 Adding a new DA method
@@
  • This class should implement a get_estimator() function, which returns a class inheriting from sklearn.BaseEstimator and accepts sample_weight as fit parameter. In the benchmark we used the Domain Adaptation toolbox SKADA (Gnassounou et al., 2024) that provides many DA estimatos with correct interface.
@@
def __init__(self, param1=10, param2=’auto’):
@@
’param1’: [10, 100],
’param2’: [’auto’, ’manual’]
@@ D.2 Adding a new dataset
# sample_domain>0 for source samples
# sample_domain<0 for target samples
@@
Appendix E Benchmark detailed results
E.1 Results per datasets
@@
These detailed tables where cell in green denote a gain wrt Train Src (average outside of standard deviation of Train Src) better illustrate the challenges of domain adaptation (DA) methods. They show that not all shifts are equivalent within a given dataset. For example, Table 12 reveals that only 3 shifts in the AmazonReview dataset present a DA problem (defined as a $>3\%$ difference in accuracy between Train Src and Train Tgt). While for the other shifts, we achieve similar performance whether we train on source or target data. Additionally, some specific shifts present a DA problem that no method can successfully address. This can be seen in the dsl → amz shift in the Office31 dataset, as shown in Table 7. Finally, some DA methods perform consistently across all shifts within a dataset, as demonstrated by the results for the 20Newsgroup dataset in Table 11.
Table 5: Accuracy score for MNIST/USPS dataset for each shift compared for all the methods. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
@@
Reweighting Dens. RW 0.66 $\pm$ 0.02 0.42 $\pm$ 0.02 0.54 $\pm$ 0.02 13.25
Disc. RW 0.6 $\pm$ 0.02 0.4 $\pm$ 0.02 0.5 $\pm$ 0.02 19.00
Gauss. RW 0.11 $\pm$ 0.01 0.11 $\pm$ 0.01 0.11 $\pm$ 0.01 20.00
KLIEP 0.66 $\pm$ 0.02 0.43 $\pm$ 0.02 0.54 $\pm$ 0.02 13.25
KMM 0.64 $\pm$ 0.02 0.41 $\pm$ 0.03 0.52 $\pm$ 0.02 18.00
NN RW 0.66 $\pm$ 0.02 0.42 $\pm$ 0.02 0.54 $\pm$ 0.02 12.00
MMDTarS 0.66 $\pm$ 0.02 0.42 $\pm$ 0.02 0.54 $\pm$ 0.02 12.75
@@
MMD-LS 0.66 $\pm$ 0.02 0.47 $\pm$ 0.02 0.56 $\pm$ 0.02 8.25
Subspace JPCA 0.66 $\pm$ 0.02 0.43 $\pm$ 0.02 0.54 $\pm$ 0.02 12.00
SA 0.71 $\pm$ 0.03 0.36 $\pm$ 0.11 0.54 $\pm$ 0.07 12.00
TCA 0.08 $\pm$ 0.07 0.11 $\pm$ 0.02 0.09 $\pm$ 0.05 21.00
TSL 0.66 $\pm$ 0.02 0.43 $\pm$ 0.02 0.54 $\pm$ 0.02 10.50
Other JDOT 0.73 $\pm$ 0.02 0.53 $\pm$ 0.02 0.63 $\pm$ 0.02 3.50
@@
Table 6: Accuracy score for MNIST/USPS dataset for each shift compared for all the Deep DA methods. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
@@
Train Src 0.94 0.76 0.85 5.0
Train Tgt 0.99 0.99 0.99 1.0
DANN 0.94 0.88 0.91 4.0
DeepCORAL 0.97 0.89 0.93 2.5
DeepJDOT 0.96 0.9 0.93 2.5
Table 7: Accuracy score for Office31 dataset for each shift compared for all the methods. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.